# Hyperdimensional Probe: Decoding LLM Representations via Vector Symbolic Architectures

### v1

@@ H.2 Applicability to other downstream tasks
Appendix I Question-answering setting from Section˜5.3
@@
Figure 13: Silhouette scores for varying numbers of clusters, computed using a random sample of 10,000 textual inputs from $\mathcal{S}$. The six language models have varied layer counts (see Table˜3), which results in different maximum possible cluster numbers.
@@
In contrast, the computational demands of our VSA-based methodology is relatively low. The most resource-intensive stage was training our neural VSA encoder, but due to its modest size (ranging from $55M$ to $71M$ parameters, see Appendix˜C), this process remained lightweight. We performed this training on a single GPU, though it could easily be handled with much less powerful and lower-memory GPUs. The probing stage is then composed of simple vector multiplications (unbinding, Section˜3), after loading the heavy LLM and our lightweight trained neural VSA encoder into memory (from 800 MB of the 55M version to 1 GB of the biggest one). Future research could explore even further reducing the latent dimension of our neural VSA encoder (Appendix˜C) or adopt VSA encodings with lower dimensionality (e.g. $D=512$, leading to a more lightweight encoder.

### v2

@@ H.2 Applicability to other downstream tasks
Appendix I Question-answering setting from Section˜6
@@
Figure 13: Silhouette scores for varying numbers of clusters, computed using a random sample of 10,000 textual inputs from $\mathcal{S}$. The six language models have varied layer counts (see Table˜4), which results in different maximum possible cluster numbers.
@@
In contrast, the computational demands of our VSA-based methodology is relatively low. The most resource-intensive stage was training our neural VSA encoder, but due to its modest size (ranging from $55M$ to $71M$ parameters, see Appendix˜C), this process remained lightweight. We performed this training on a single GPU, though it could easily be handled with much less powerful and lower-memory GPUs. Regarding GPU usage, we trained our encoder for approximately 8 hours on each LLM’s embeddings, though the process could have been shortened with a less conservative early stopping criterion or by reducing the amount of training data.
The probing stage is then composed of simple vector multiplications (unbinding, Section˜3), after loading the heavy LLM and our lightweight trained neural VSA encoder into memory (from 800 MB of the 55M version to 1 GB of the biggest one). Future research could explore even further reducing the latent dimension of our neural VSA encoder (Appendix˜C) or adopt VSA encodings with lower dimensionality (e.g. $D=512$, leading to a more lightweight encoder. The time cost of probing is thus comparable to simple LLM inference with a slight increase due to feedforward our lightweight trained model, with the time demand for vector multiplications being negligible. Accordingly, the GPU hours for probing depends mainly on the amount of test data and model size; for example, we used around 95 hours of GPU computation for probing embeddings of Llama 3.1-8B on $\bar{\mathcal{S}}\approx 10^{5}$, processing each instance in around 3 seconds.